solve for x and y
5x - 2y = 12
4x + 3y = 8

Answer by mananth(16946)

5x - 2y = 12 ---1
4x + 3y = 8 ---2
Multiply eq-- 1 by 3
& eq---2 by 2
we get
15x -6y = 36
8x + 6y = 16
Add the two equations
we get
23x = 52
x= 52/23
Put the value of x in eq. 1
15 * ( 52/23) -2y = 12
-2y= (12 *23)/(52*15)
y= - (12*23)/2*52*15)
Simplify the fraction to get the value of y